OpenStudy (anonymous):

Which of the following graphs best represents the solution to the pair of equations below? y = −x + 6 y = 2x − 3

OpenStudy (anonymous):

A coordinate plane is shown with two lines graphed. One line passes through the y axis at 6 and through the x axis at 6. The other line passes through the y axis at negative 3 and the x axis at 1.5. The lines intersect at the point 3 comma 3.

OpenStudy (anonymous):

A coordinate plane is shown with two lines graphed. One line passes through the y axis at 6 and the x axis at 6. The other line passes through the y axis at 3 and the x axis at negative 1.5. The lines intersect at 1 comma 5.

OpenStudy (anonymous):

A coordinate plane is shown with two lines graphed. One line passes through the y axis at 6 and the x axis at negative 6. The other line passes through the y axis at negative 3 and the x axis at negative 1.5. The lines intersect at negative 3 and 3.

OpenStudy (anonymous):

A coordinate plane is shown with two lines graphed. One line passes through the y axis at 6 and the x axis at negative 6. The other line passes through the y axis at 3 and the x axis at 1.5. The lines intersect at negative 1 comma 5.

OpenStudy (anonymous):

@Mackenzie_Willa

OpenStudy (anonymous):

@peachpi

OpenStudy (lynfran):

to find where the graph intersect we equate the two equations so -x+6=2x-3 -x-2x=-3-6 -3x=-9 x=3 to find y when x=3 we sub that into any 1 of the give equation to get.. y=-(3)+6 y=3